Write a Quadratic Function having the given solutions:6,-2.
Answer by ankor@dixie-net.com(22740)

Write a Quadratic Function having the given solutions:6,-2.
:
These solutions result from factors:
(x-6)(x+2) = 0
FOIL
x^2 + 2x - 6x - 12 = 0
f(x) = x^2 - 4x - 12